OpenStudy (xxlyh):

can someone tell me if my answer was right? An average clementine is shaped like a sphere and has a diameter of 3.5 inches. What is the volume of the fruit? Use 3.14 for pi. Round your answer to the nearest hundredth. (4 points) 20.45 cubic inches 21.34 cubic inches 22.44 cubic inches 23.78 cubic inches

OpenStudy (xxlyh):

I chose C

OpenStudy (md152727):

\(V=\dfrac{4}{3}\pi r^2\)

OpenStudy (md152727):

d=3.5 r=d/2 r=3.5/2

OpenStudy (md152727):

Does this help?

OpenStudy (xxlyh):

r^3 * 4 * 3.14 /3 r = 1.75

OpenStudy (xxlyh):

I used that and got 22.43

OpenStudy (md152727):

or C Please close this question